# A note on the electrostatic Born–Infeld equation with radial charge density

### 06 Apr 2026 v3

@@ Abstract
In this note, we will give a new proof of the solvability of the electrostatic Born–Infeld equation with radial charge by using the conformal method and the Spacetime Positive Energy Theorem.
@@ 1 Introduction
The Born–Infeld Lagrangian plays a fundamental role in field theory. It was introduced to address the violation of the Principle of Finite Energy in the classical electrostatic Maxwell model. In the vacuum case, it leads to the equation
@@
where $\rho$ is the charge density and $u$ represents the electric potential. This equation is called the electrostatic Born–Infeld equation and has attracted significant attention in recent years. A classical and widely used approach to solving this problem is the variational method. In this framework, the equation arises as the Euler–Lagrange equation of the functional
@@
and weak solutions $u$ are obtained by looking for minimizers of $I$. However, most existing results rely on restrictive assumptions on $\rho$, such as radial symmetry or smallness conditions (see [2, 3]).
In this work, we revisit the case of radial charge densities and propose a new proof of existence in view of general relativity. The key observation is that, in classical relativity, the equation (1.1a) is exactly the mean curvature equation in Lorentz–Minkowski spacetime $\mathbb{L}^{n+1}$. Indeed, any spacelike hypersurface of $\mathbb{L}^{n+1}$ can be written as a graph $\Sigma=\{(x,u(x))~|~x\in\mathbb{R}^{n}\}$, and if $\rho$ denotes the mean curvature of its second fundamental form, then the pair $(\rho,u)$ automatically satisfies (1.1a). Moreover, since the induced metric $h$ of $\Sigma$ in this case is expressed by
@@
if $(\Sigma,h)$ is sufficiently asymptotically flat so that $h-\delta_{\text{Euc}}=O(r^{-p})$ at infinity for some $p>2$, it follows that the graph function $u$ converges to a constant at infinity, which we can normalize to be zero, and so $u$ also holds the decay condition (1.1b).
This geometric insight proposes us a different approach that instead of solving (1.1) directly, we will aim to construct an asymptotically flat (AF) spacelike hypersurface of $\mathbb{L}^{n+1}$ with prescribed mean curvature $\rho$. However, how to achieve so without explicitly using (1.1) is a challenge.
To deal with this, we will rely on two tools from general relativity: the conformal method and the Spacetime Positive Energy Theorem (PET). More precisely, it is well-known by the rigidity part of the Spacetime PET that if an AF manifold $(M,g)$, together with a tensor $k$, forms a vacuum initial data set for the Cauchy problem in general relativity and has zero ADM mass, then $(M,g)$ is isometric to a spacelike hypersurface of $\mathbb{L}^{n+1}$ with second fundamental form $k$. Therefore, in order to construct the desired hypersurface, it suffices to find a vacuum AF initial data set $(M,g,k)$ such that $\text{tr}_{g}k=\rho$ and the ADM mass of $(M,g)$ vanishes. For this task, it is natural to think of using the conformal method, a classical approach allowing one to construct initial data sets from scratch with prescribed mean curvature. This approach is feasible, as we will see below, at least when $\rho$ is radial, the construction becomes quite simple. Working in the weighted Hölder spaces introduced in Subsection 2.2, we obtain the following result.
@@
Then there exists an asymptotically flat spake-like hypersurface $(\Sigma,h)$ of the Lorentz-Minkowski spacetime $\mathbb{L}^{n+1}$ with mean curvature $\rho$ and $h-\delta_{\text{Euc}}\in C^{3,\alpha}_{-2q+2}$. In particular, the electrostatic Born-Infeld equation (1.1) admits at least one solution $u$.
In comparison with the result in [2], our result is a bit weaker, since we require regularity assumptions on $\rho$ in $\mathbb{R}^{n}$. Nevertheless, the approach has a clear advantage: the solutions obtained in Theorem 1.1 are classical, whereas those derived via the variational method are only weak solutions. Moreover, our method avoids a difficulty in the variational approach, namely, showing that a minimizer of $I$ actually solves (1.1).
It is also worth noting that, as mentioned in Section 4, the theorem can be further strengthened by assuming only that $q>2$. However, we prefer the stated version since it seems to allow a generalization to a larger class of $\rho$, rather than being restricted to the radial case.
The outline of this article is as follows. Section 2 contains a brief summary of the conformal method and the Spacetime PET. In Section 3, we give the proof of Theorem 1.1 based on these tools. In Section 4, we make some important remarks for the further discussion and potential future developments of this approach.
@@ 2 Preliminaries
In this section, we review some standard facts on the conformal method, weighted Hölder spaces and the Spacetime PET. While these notions are well defined on general manifolds, for our purposes we only restrict our attention to the Euclidean space $(\mathbb{R}^{n},\delta_{\text{Euc}})$. For a treatment of the general case, we refer the reader to [1, 6].
@@ 2.1 Conformal method
An asymptotically flat (AF) manifold $(M,g)$ of $n$ dimensions, with $n\geq 3$, coupled with a symmetric $(0,2)$-tensor $k$, is called a vacuum initial data set for the Cauchy problem in general relativity if $(M,g,k)$ satisfies the system
@@
where $R_{g}$ is the scalar curvature of $g$. These equations are called the vacuum Einstein constraint equations and the study of (2.1) is a topical issue. Using the conformal method, to construct solutions to (2.2) starting from the Euclidean space, let $\tau$ be a scalar function and let $\sigma$ be a trace-free and divergence–free symmetric $(0,2)$-tensor on $(\mathbb{R}^{n},\delta_{\text{Euc}})$, one is required to find a positive function $\varphi$ tending to $1$ at infinity and a 11-form $W$ satisfying
@@
is a solution to the vacuum constraint equations (2.1). In this situation, we remark that $\text{tr}_{g}k=\rho$, therefore, $\rho$ is called the mean curvature in free data of this method.
@@ 2.3 Spacetime Positive Energy Theorem
Recall that the following definition and results are formulated for a general AF manifold. Here, however, we restrict our discussion to $\mathbb{R}^{n}$. Similarly, for simplicity and in line with our purposes, we consider the PET only in the vacuum case.
@@
Bartnik in [1] showed that under the decay condition (2.6), the mass is a geometric invariant. A long-standing conjecture in general relativity states that the ADM mass of a vacuum AF initial data set is positive unless it is a space-like hypersurface of the Lorentz–Minkowski spacetime. This conjecture was proven to be true under suitable decay assumptions of $(g,k)$ at infinity and structural conditions on the manifold. The two most well-known results for this problem are the proof by Chruściel–Maerten [5] for spin manifolds in arbitrary dimensions, and the proof by Eichmair [7] for general manifolds in dimensions less than eight. We note that the decay rate assumptions on $(g,k)$ at infinity in these results differ slightly. Since $\mathbb{R}^{n}$ is spin for all $n\geq 3$, in the statement below we adopt the stronger of the two sets of assumptions.
@@
We are now ready to construct spacelike hypersurfaces of the Lorentz–Minkowski spacetime $\mathbb{L}^{n+1}$ with a prescribed radial mean curvature. The construction will be carried out in three steps, corresponding to the subsections of this section. First, since the approach relies on the conformal method, we study how radial solutions of the conformal equations behave when the mean curvature $\rho$ is radial. Next, using this analysis, we give a detailed construction of an asymptotically flat solution to the vacuum constraints that has the prescribed radial mean curvature $\rho$. Finally, we show that this solution has zero mass, and hence, by the rigidity part of the Spacetime PET, it is isometric to a spacelike hypersurface of $\mathbb{L}^{n+1}$, which is our object.
@@ 3.1 Behavior of radial solutions to the conformal equation
Here and subsequently, $r$ is the usual Euclidean distance and we denote by $f^{\prime}$ the derivative of $f$ with respect to $r$. Four our purpose, we only restrict our attention to the set of all radial solutions to (3.1). The following result plays a key role in our analysis.
@@
Differentiating (3.3) with respect to $i$ and summing all equations of the new system, we obtain $\Delta\Big(2\sum_{j=1}^{n}\partial_{j}W_{j}\Big)=\Delta f.$ Therefore, by Proposition 2.2(a), we have $2\sum_{j=1}^{n}\partial_{j}W_{j}=f.$ Taking into account (3.3), we get $\Delta W_{i}={1\over 2}\partial_{i}f$. Since $f$ is radial, it follows by computations that
@@
To simplify the formula, it is not difficult to check that
@@
Step 3. $\varphi$ is increasing. We remind the reader that $\varphi^{\prime}$ was assumed to be different from 0 a.e in $\mathbb{R}^{n}$ for simplicity. In view of (3.9)–(3.11), we see that the necessary condition for $(\varphi,\rho)$ to be a solution to (3.1) is
@@
We will show that this condition is equivalent to the fact that $\varphi$ is increasing. In fact, if $\varphi^{\prime}>0$ a.e in $\mathbb{R}^{n}$, (3.12) is obvious. Conversely, assume that (3.12) holds. Since $\varphi(0)>0$, we have $2n\varphi+Nr\varphi^{\prime}>0$ near 0. It then follows by (3.12) that $\varphi^{\prime}\geq 0$ near 0. Therefore, if $\varphi^{\prime}<0$ somewhere, then there exists a convergent sequence $\{r_{m}\}$ such that $\varphi^{\prime}(r_{m})<0$ and $\varphi^{\prime}(r_{m})\to 0$. This leads to the contradiction that $0\leq\varphi^{\prime}(r_{m})\big(2n\varphi(r_{m})+Nr_{m}\varphi^{\prime}(r_{m})\big)<0$. Therefore, we have $\varphi^{\prime}>0$ a.e, and so the inequality (3.12) holds as we assumed in (3.10). The proof is completed. ∎
@@ 3.2 Existence of constraint solutions with prescribed radial mean curvature
$\big(g_{ij},k_{ij}\big):=\bigg(\varphi^{N-2}\delta_{ij},\,{\tau\over n}\varphi^{N-2}\delta_{ij}-\varphi^{-2}\Big({\delta_{ij}\over nr^{n}}-{x_{i}x_{j}\over r^{n+2}}\Big)\int_{0}^{r}s^{n}\varphi^{N}\rho^{\prime}\,ds\bigg),$ (3.13)
@@
We note that, by substituting the expression (3.4) for $LW$ into (2.4), the conformal method reduces the problem to showing that the system (3.1) admit a radial positive solution $\varphi$ with $\varphi-1\in C^{3,\alpha}_{-2\beta+2}(\mathbb{R}^{n})$.
@@
- If $T(t,\phi)=\phi$ with $(t,\phi)\in(0,1]\times RL^{\infty}$, then $\phi/t$ is a radial solution to the conformal equations (3.1) corresponding to the seed data $(\delta_{\text{Euc}},t^{N}\rho)$. By Proposition 3.1, it follows that $\phi/t$ is increasing, and hence $\|\phi/t\|_{L^{\infty}}=1$. In particular, the set $K=\big\{(t,\phi)\in(0,1]\times RL^{\infty}~\big|~T(t,\phi)=\phi\big\}$ is bounded.
From these observations, once $T$ is proven to be continuous and compact in $0,1]\times RL^{\infty}$, the Leray–Schauder fixed point ensures that $T(1,.)$ has a fixed point in $RL^{\infty}$ which is exactly what we have desired. Moreover, in view of (3.17), since $\mathcal{V}$ is continuous and $\mathcal{I}$ is continuous and compact, it follows that $T$ will be continuous and compact once we establish the continuity of $\mathcal{L}$. Therefore, it remains to prove that $\mathcal{L}$ is a continuous operator. The argument we give here is essentially the same as in [9, 10], which give the corresponding result for compact manifolds.
@@
It follows that $F_{\psi}\in C\big([0,1]\times C^{2,\alpha}_{-\beta+1},L(C^{3,\alpha}_{-2\beta+2},C^{1,\alpha}_{-2\beta})\big)$, where we denote $\big(L(C^{3,\alpha}_{-2\beta+2},C^{1,\alpha}_{-2\beta})\big)$ the Banach space of all linear continuous maps from $C^{3,\alpha}_{-2\beta+2}$ into $C^{1,\alpha}_{-2\beta}$. In particular, setting $\psi_{0}=\mathcal{L}(t,W)$ we have

### 16 Apr 2026 v5

@@ Abstract
In this note, we present a new proof of the solvability of the electrostatic Born–Infeld equation with radial charge, based on the conformal method and the Spacetime Positive Energy Theorem. An advantage of this approach is that the resulting solutions are automatically classical and spacelike. 2020 Mathematics Subject Classification. 35J93 (Primary), 35Q60, 53C21, 83C05 (Secondary).Key words and phrases. Born-Infeld equation, mean curvature equation, conformal method, Spacetime Positive Energy Theorem.
@@ 1 Introduction
The Born–Infeld Lagrangian is a foundation of field theory, introduced to deal with the failure of classical electrostatic Maxwell theory to satisfy the Principle of Finite Energy. In the vacuum case, it leads to the relation
@@
where $\rho$ is the charge density and $u$ represents the electric potential. This equation is called the electrostatic Born–Infeld equation and has received significant attention in recent years. The most common approach to solving this problem is the variational method, in which the equation is interpreted as the Euler–Lagrange equation of the functional
@@
and one looks for a weak solution $u$ to (1.1) as the minimizer of $I$. Until now, there have been two main existence results obtained for (1.1) under assumptions of $\rho$: the first one given in [2] requires that $\rho$ is radially distributed, while the second one [3] relies on a smallness condition on $\rho$.
In this work, we are interested in the case of radial charge densities and provide a new proof of existence in view of general relativity. The key observation is that, in classical relativity, the equation (1.1a) is exactly the mean curvature equation in Lorentz–Minkowski spacetime $\mathbb{L}^{n+1}$. Indeed, any spacelike hypersurface of $\mathbb{L}^{n+1}$ can be written as a graph $\Sigma=\{(x,u(x))~|~x\in\mathbb{R}^{n}\}$, and if $\rho$ denotes the mean curvature of its second fundamental form, then the pair $(\rho,u)$ automatically satisfies (1.1a). In addition, since the induced metric $h$ of $\Sigma$ in this case is expressed by
@@
if $(\Sigma,h)$ is sufficiently asymptotically flat so that $h-\delta_{\text{Euc}}=O(r^{-p})$ at infinity for some $p>2$, it follows that the graph function $u$ converges to a constant at infinity, which we can normalize to zero, and so $u$ also holds (1.1b).
The geometric insight proposes a different approach that instead of solving (1.1) directly, we will aim to construct an asymptotically flat (AF) spacelike hypersurface of $\mathbb{L}^{n+1}$ with prescribed mean curvature $\rho$.
[…]
Figure 1: Variation method and an Approach based on general relativity
To deal with this, we rely on two tools from general relativity: the conformal method and the Spacetime Positive Energy Theorem (PET). More precisely, as illustrated in Figure 1, which presents a schematic of the two methods discussed, it is well-known by the rigidity part of the Spacetime PET that if an AF manifold $(M,g)$, together with a tensor $k$, forms a vacuum initial data set for the Cauchy problem in general relativity and has zero ADM mass, then $(M,g)$ is isometric to a spacelike hypersurface of $\mathbb{L}^{n+1}$ with second fundamental form $k$. Therefore, to construct the desired hypersurface, we only need to find a vacuum AF initial data set $(M,g,k)$ such that $\text{tr}_{g}k=\rho$ and its mass is zero. For this task, it is natural to think of using the conformal method, a traditional to construct initial data sets from scratch with prescribed mean curvature. When $\rho$ is radial, as we will see below, this approach yields a feasible and quite simple construction. More precisely, working within the weighted Hölder spaces introduced in Subsection 2.2, we obtain the following result.
@@
Then there exists an asymptotically flat spacelike hypersurface $(\Sigma,h)$ of the Lorentz-Minkowski spacetime $\mathbb{L}^{n+1}$ such that $h-\delta_{\text{Euc}}\in C^{3,\alpha}_{-2q+2}$ and its mean curvature equals to $\rho$. In particular, the electrostatic Born-Infeld equation (1.1) admits at least one spacelike solution $u\in C^{2,\alpha}_{-q+2}$.
In comparison with the result in [2], our result is slightly weaker, since we require Hölder regularity for $\rho$ in $\mathbb{R}^{n}$. Nevertheless, the approach has an advantage that the resulting solutions in Theorem 1.1 are spacelike and classical, whereas those derived via the variational method are only weak solutions. Moreover, our method avoids the difficulty in the variational approach, namely, showing that a minimizer of $I$ solves (1.1).
It is also worth noting that, as mentioned in Section 3, Theorem 1.1 can be further strengthened by assuming only that $q>2$. However, we prefer the stated version since it seems to allow a generalization to a larger class of $\rho$, rather than being restricted to the radial case.
The outline of the article is as follows. In Section 2, we give a brief summary of the conformal method and the Spacetime PET. In Section 3, the proof of Theorem 1.1 based on these tools is established.
@@ 2 Preliminaries
In this section, we review some standard facts on the conformal method, weighted Hölder spaces and the Spacetime PET. While these notions are well defined on general manifolds, for our purposes we restrict our attention to the Euclidean space $(\mathbb{R}^{n},\delta_{\text{Euc}})$. For a treatment of the general case, we refer the reader to [1, 6].
@@ 2.1 Conformal method
An AF manifold $(M,g)$ of $n$ dimensions, with $n\geq 3$, coupled with a symmetric $(0,2)$-tensor $k$, is called a vacuum initial data set for the Cauchy problem in general relativity if $(M,g,k)$ satisfies the system
@@
where $R_{g}$ is the scalar curvature of $g$. These equations are called the vacuum Einstein constraint equations and the study of (2.1) is a topical issue. Using the conformal method, to construct solutions to (2.1) grounded in Euclidean space, let $\rho$ be a scalar function and let $\sigma$ be a trace-free and divergence–free symmetric $(0,2)$-tensor on $(\mathbb{R}^{n},\delta_{\text{Euc}})$, one is required to find a positive function $\varphi$ tending to $1$ at infinity and a 11-form $W$ satisfying
@@
is a solution to the vacuum constraint equations (2.1). In this case, we remark that $\text{tr}_{g}k=\rho$, therefore, $\rho$ is called the mean curvature in free data of this method.
@@ 2.3 Spacetime Positive Energy Theorem
Recall that the following definition and results are formulated for a general AF manifold. Here, however, we restrict our discussion to $\mathbb{R}^{n}$. Similarly, for simplicity and in line with our purposes, we consider the theorem only in the vacuum case.
@@
Bartnik in [1] showed that under the decay condition (2.6), the mass is a geometric invariant. A long-standing conjecture in general relativity states that the ADM mass of a vacuum AF initial data set is positive unless it is a spacelike hypersurface of the Lorentz–Minkowski spacetime. This conjecture was proven to be true under suitable decay assumptions of $(g,k)$ at infinity and structural conditions on the manifold. The two most well-known results for this problem are the proof by Chruściel–Maerten [5] for spin manifolds in arbitrary dimensions, and the proof by Eichmair [7] for general manifolds in dimensions less than eight. We remark that the decay rate assumptions on $(g,k)$ at infinity in these results differ slightly. However, since our considered manifold $\mathbb{R}^{n}$ is spin for all $n\geq 3$, in the statement below we adopt the stronger of the two sets of decay rate assumptions.
@@
We are now ready to construct spacelike hypersurfaces of the Lorentz–Minkowski spacetime $\mathbb{L}^{n+1}$ with a prescribed radial mean curvature. The construction will be carried out in three steps, corresponding to the subsections of this section. First, since the approach relies on the conformal method, we study how radial solutions of the conformal equations behave when the mean curvature $\rho$ is radial. Next, using this analysis, we give a detailed construction of an AF solution to the vacuum constraints that has the prescribed radial mean curvature $\rho$. Finally, we show that this solution has zero mass, and hence, by the rigidity part of the Spacetime PET, it is isometric to a spacelike hypersurface of $\mathbb{L}^{n+1}$, which is our object.
@@ 3.1 Behavior of radial solutions to the conformal equation
Here and subsequently, $r$ is the usual Euclidean distance and we denote by $f^{\prime}$ the derivative of $f$ with respect to $r$. For our purpose, we restrict our attention to the set of all radial solutions to (3.1). The following result plays a key role in our analysis.
@@
Differentiating (3.3) with respect to $i$ and summing all equations of the new system, we obtain $\Delta\Big(2\sum_{j=1}^{n}\partial_{j}W_{j}\Big)=\Delta f.$ Therefore, by Proposition 2.2(a), we have $2\sum_{j=1}^{n}\partial_{j}W_{j}=f.$ Taking into account (3.3), we get $\Delta W_{i}={1\over 2}\partial_{i}f$. Since $f$ is radial, a direct computation shows that
@@
To simplify the formula, one can verify that
@@
Step 3. $\varphi$ is increasing. Recall that $\varphi^{\prime}$ was assumed to be different from 0 a.e in $\mathbb{R}^{n}$ for simplicity. In view of (3.9)–(3.11), we see that the necessary condition for $(\varphi,\rho)$ to be a solution to (3.1) is
@@
We will show that this condition is equivalent to the fact that $\varphi$ is increasing. In fact, if $\varphi^{\prime}>0$ a.e in $\mathbb{R}^{n}$, (3.12) is obvious. Conversely, assume that (3.12) holds. Since $\varphi(0)>0$, we have $2n\varphi+Nr\varphi^{\prime}>0$ near 0. It then follows by (3.12) that $\varphi^{\prime}\geq 0$ near 0. Therefore, if $\varphi^{\prime}<0$ somewhere, then there exists a convergent sequence $\{r_{m}\}$ such that $\varphi^{\prime}(r_{m})<0$ and $\varphi^{\prime}(r_{m})\to 0$. This yields the contradiction $0\leq\varphi^{\prime}(r_{m})\big(2n\varphi(r_{m})+Nr_{m}\varphi^{\prime}(r_{m})\big)<0$. Therefore, we have $\varphi^{\prime}>0$ a.e, and so the inequality (3.12) holds as we assumed in (3.10). The proof is completed. ∎
@@ 3.2 Existence of constraint solutions with prescribed radial mean curvature
$\big(g_{ij},k_{ij}\big):=\bigg(\varphi^{N-2}\delta_{ij},\,{\rho\over n}\varphi^{N-2}\delta_{ij}-\varphi^{-2}\Big({\delta_{ij}\over nr^{n}}-{x_{i}x_{j}\over r^{n+2}}\Big)\int_{0}^{r}s^{n}\varphi^{N}\rho^{\prime}\,ds\bigg),$ (3.13)
@@
We note that, by substituting the expression (3.4) for $LW$ into (2.4), it follows from the conformal method that the setting $(g,k)$ given in (3.13) is automatically a solution to the constraint once $(\varphi,\rho)$ solves the conformal equations (3.1). Therefore, it suffices to prove that (3.1) admit a radial positive solution $\varphi$ with $\varphi-1\in C^{3,\alpha}_{-2\beta+2}(\mathbb{R}^{n})$.
@@
- If $T(t,\phi)=\phi$ with $(t,\phi)\in(0,1]\times RL^{\infty}$, then $\phi/t$ is a radial solution to the conformal equations (3.1) associated with the seed data $(\delta_{\text{Euc}},t^{N}\rho)$. By Proposition 3.1, it follows that $\phi/t$ is increasing, and hence $\|\phi/t\|_{L^{\infty}}=1$. In particular, the set $K=\big\{(t,\phi)\in(0,1]\times RL^{\infty}~\big|~T(t,\phi)=\phi\big\}$ is bounded.
From these observations, once $T$ is proven to be continuous and compact in $0,1]\times RL^{\infty}$, the Leray–Schauder fixed point ensures that $T(1,.)$ has a fixed point in $RL^{\infty}$ which is exactly what we desire. Moreover, in view of (3.17), since $\mathcal{V}$ is continuous and $\mathcal{I}$ is continuous and compact, it follows that $T$ will be continuous and compact once we establish the continuity of $\mathcal{L}$. Therefore, it remains to prove that $\mathcal{L}$ is a continuous operator. The argument we give here is essentially the same as in [9, 10], which give the corresponding result for compact manifolds.
@@
It follows that $F_{\psi}\in C\big([0,1]\times C^{2,\alpha}_{-\beta+1},L(C^{3,\alpha}_{-2\beta+2},C^{1,\alpha}_{-2\beta})\big)$, where we denote $L(C^{3,\alpha}_{-2\beta+2},C^{1,\alpha}_{-2\beta})$ the Banach space of all linear continuous maps from $C^{3,\alpha}_{-2\beta+2}$ into $C^{1,\alpha}_{-2\beta}$. In particular, setting $\psi_{0}=\mathcal{L}(t,W)$ we have